# Unlocking the Potential of Model Calibration in Federated Learning

### v2

@@ B.2 Additional Results of Calibration on FL
Calibration Method FedAvg McMahan et al. [2017] FedProx Sahu et al. [2018] Scaffold Karimireddy et al. [2019] FedDyn Acar et al. [2021] FedNova Wang et al. [2020b]
Uncal. 79.23 9.60 2.13 78.52 7.41 2.52 78.27 15.91 5.45 77.65 10.11 8.17 79.06 9.88 2.82
Focal Lin et al. [2017] 78.46 4.79 1.51 78.37 11.74 2.96 78.78 12.36 5.69 73.91 6.12 5.49 78.37 11.61 4.40
LS Müller et al. [2019] 78.78 4.30 1.73 78.08 13.38 3.83 78.25 14.49 5.19 76.49 8.20 10.07 78.63 10.37 4.00
BS Brier [1950] 74.90 5.27 1.50 72.97 8.82 3.03 76.54 9.32 6.06 77.41 15.87 9.85 76.06 5.04 2.01
MMCE Kumar et al. [2018] 78.49 4.82 1.23 78.54 7.69 2.33 78.26 14.61 5.21 75.03 8.60 6.30 78.22 8.02 3.01
FLSD Mukhoti et al. [2020] 76.76 8.52 2.04 78.01 15.46 3.73 77.26 14.43 5.36 76.09 8.60 6.30 76.50 8.54 2.03
MbLS Liu et al. [2021] 78.73 9.17 2.00 78.30 9.01 2.75 78.57 16.01 5.06 76.65 8.09 8.28 79.47 8.63 1.94
DCA Liang et al. [2020] 79.33 5.37 1.71 78.51 6.35 2.16 78.23 12.02 5.09 77.40 7.81 5.46 79.39 8.56 2.19
MDCA Hebbalaguppe et al. [2022b] 79.40 7.04 2.03 78.52 8.84 3.67 77.41 12.33 6.04 77.12 8.16 7.18 79.42 7.73 2.00
NUCFL (DCA+COS) 79.29 5.09 1.35 78.51 5.26 2.09 78.22 10.19 4.95 77.54 6.00 4.60 79.84 6.76 1.90
NUCFL (MDCA+COS) 79.60 6.79 1.99 78.51 5.67 2.15 78.82 10.97 5.02 77.69 6.10 4.74 79.15 6.91 2.02
NUCFL (DCA+L-CKA) 79.50 5.85 2.08 79.61 5.34 1.93 78.35 11.35 5.08 77.18 6.03 4.07 79.57 6.69 1.99
NUCFL (MDCA+L-CKA) 79.27 5.48 2.08 79.61 5.69 2.06 78.32 11.25 5.08 76.81 5.76 4.29 79.26 7.03 2.03
NUCFL (DCA+RBF-CKA) 79.42 5.59 2.01 79.61 5.44 1.83 78.59 9.05 4.73 76.66 5.91 4.19 79.49 6.70 2.03
NUCFL (MDCA+RBF-CKA) 79.87 5.68 2.04 79.60 5.57 2.03 78.59 10.06 4.97 76.88 5.44 4.69 79.46 7.15 2.02
@@
Table 11: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under IID scenario on the CIFAR-10 dataset. Values in boldface indicate the best calibration across all methods, achieved without compromising accuracy compared to the uncalibrated model.
@@
Calibration Method FedAvg McMahan et al. [2017] FedProx Sahu et al. [2018] Scaffold Karimireddy et al. [2019] FedDyn Acar et al. [2021] FedNova Wang et al. [2020b]
Uncal. 75.33 9.69 2.31 74.87 5.36 2.84 75.19 7.99 2.98 74.00 18.84 5.88 75.37 10.67 2.34
Focal Lin et al. [2017] 73.60 5.04 1.63 74.92 9.71 3.16 74.04 9.44 3.04 71.26 16.35 5.27 74.07 5.40 3.70
LS Müller et al. [2019] 75.00 3.87 2.04 74.84 11.26 4.21 75.90 8.38 2.47 74.63 15.35 4.37 75.50 5.80 2.68
BS Brier [1950] 71.22 5.40 1.76 73.89 9.27 2.71 73.03 6.03 2.20 73.22 16.23 5.33 71.70 6.34 2.41
MMCE Kumar et al. [2018] 74.65 9.08 3.72 75.07 9.29 2.81 75.16 6.37 2.95 71.34 14.53 5.17 74.16 6.82 3.65
FLSD Mukhoti et al. [2020] 70.74 8.81 2.62 74.52 13.24 4.24 73.50 6.37 2.63 73.39 18.96 6.78 70.75 8.21 2.37
MbLS Liu et al. [2021] 75.57 8.70 2.10 74.96 6.31 2.63 75.68 8.67 3.96 74.16 16.62 5.90 75.59 9.03 2.63
DCA Liang et al. [2020] 75.34 5.45 1.67 74.99 4.14 2.42 75.12 6.23 2.55 73.88 10.00 4.86 74.43 6.22 2.32
MDCA Hebbalaguppe et al. [2022b] 75.42 6.44 2.25 74.92 6.86 2.68 74.16 6.20 3.07 75.70 9.27 4.57 75.05 6.49 2.24
NUCFL (DCA+COS) 75.39 5.10 1.42 74.93 3.51 2.13 75.13 5.54 1.76 74.98 6.17 4.14 75.91 5.74 2.22
NUCFL (MDCA+COS) 75.47 6.03 2.09 74.97 3.47 2.08 75.36 5.60 1.71 73.10 5.03 4.06 75.33 5.54 2.03
NUCFL (DCA+L-CKA) 76.96 4.76 1.22 75.76 3.66 2.15 75.34 5.23 1.41 74.68 5.42 4.24 75.29 4.99 1.70
NUCFL (MDCA+L-CKA) 76.34 4.76 1.22 75.74 3.27 1.91 75.57 4.17 1.30 74.67 5.27 4.13 76.85 6.01 2.24
NUCFL (DCA+RBF-CKA) 76.74 4.43 1.13 75.70 3.70 2.00 76.11 4.91 1.76 74.75 5.38 4.06 76.74 5.60 1.91
NUCFL (MDCA+RBF-CKA) 76.88 4.95 1.25 75.73 3.81 2.38 76.82 4.94 1.57 74.41 5.28 4.14 76.27 6.29 2.26
@@
Table 12: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under non-IID ($\alpha=0.5$) scenario on the CIFAR-10 dataset. Values in boldface indicate the best calibration across all methods, achieved without compromising accuracy compared to the uncalibrated model.

### v3

@@ B.2 Additional Results of Calibration on FL
Calibration Method FedAvg (McMahan et al., 2017) FedProx (Sahu et al., 2018) Scaffold (Karimireddy et al., 2019)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 92.48 $\pm$ 0.23 4.10 $\pm$ 0.15 1.66 $\pm$ 0.08 93.07 $\pm$ 0.27 4.49 $\pm$ 0.16 1.77 $\pm$ 0.07 91.77 $\pm$ 0.32 4.28 $\pm$ 0.19 1.66 $\pm$ 0.10
Focal (Lin et al., 2017) 91.06 $\pm$ 0.28 4.73 $\pm$ 0.17 1.98 $\pm$ 0.11 91.51 $\pm$ 0.44 5.18 $\pm$ 0.14 2.55 $\pm$ 0.07 90.85 $\pm$ 0.24 5.64 $\pm$ 0.04 2.73 $\pm$ 0.05
LS (Müller et al., 2019) 90.39 $\pm$ 0.32 3.98 $\pm$ 0.17 1.65 $\pm$ 0.19 92.96 $\pm$ 0.26 6.33 $\pm$ 0.22 3.16 $\pm$ 0.17 90.61 $\pm$ 0.66 4.22 $\pm$ 0.61 1.66 $\pm$ 0.29
BS (Brier, 1950) 92.11 $\pm$ 0.31 5.11 $\pm$ 0.30 2.55 $\pm$ 0.29 93.01 $\pm$ 0.32 6.28 $\pm$ 0.32 3.17 $\pm$ 0.19 90.00 $\pm$ 0.19 6.08 $\pm$ 0.44 3.00 $\pm$ 0.29
MMCE (Kumar et al., 2018) 92.15 $\pm$ 0.30 4.98 $\pm$ 0.29 2.52 $\pm$ 0.17 91.16 $\pm$ 0.38 3.89 $\pm$ 0.29 1.65 $\pm$ 0.11 92.75 $\pm$ 0.27 6.90 $\pm$ 0.19 3.52 $\pm$ 0.11
FLSD (Mukhoti et al., 2020) 91.47 $\pm$ 0.33 3.26 $\pm$ 0.29 1.44 $\pm$ 0.05 92.23 $\pm$ 0.28 5.93 $\pm$ 0.19 2.78 $\pm$ 0.10 90.41 $\pm$ 0.30 6.31 $\pm$ 0.18 3.17 $\pm$ 0.11
MbLS (Liu et al., 2021) 92.11 $\pm$ 0.25 4.61 $\pm$ 0.21 1.70 $\pm$ 0.19 92.00 $\pm$ 0.29 4.98 $\pm$ 0.32 2.52 $\pm$ 0.19 90.55 $\pm$ 0.44 4.30 $\pm$ 0.32 1.68 $\pm$ 0.08
DCA (Liang et al., 2020) 92.54 $\pm$ 0.23 4.00 $\pm$ 0.14 1.63 $\pm$ 0.09 93.40 $\pm$ 0.25 4.28 $\pm$ 0.15 1.65 $\pm$ 0.11 91.45 $\pm$ 0.24 4.22 $\pm$ 0.23 1.64 $\pm$ 0.10
NUCFL (DCA+COS) 92.67 $\pm$ 0.21 3.55 $\pm$ 0.11 1.51 $\pm$ 0.06 93.16 $\pm$ 0.22 3.81 $\pm$ 0.13 1.60 $\pm$ 0.09 92.10 $\pm$ 0.24 4.11 $\pm$ 0.22 1.60 $\pm$ 0.14
NUCFL (DCA+L-CKA) 92.39 $\pm$ 0.22 3.49 $\pm$ 0.20 1.46 $\pm$ 0.11 93.06 $\pm$ 0.23 3.70 $\pm$ 0.14 1.55 $\pm$ 0.11 92.73 $\pm$ 0.24 4.16 $\pm$ 0.15 1.61 $\pm$ 0.13
NUCFL (DCA+RBF-CKA) 92.58 $\pm$ 0.11 3.50 $\pm$ 0.17 1.49 $\pm$ 0.11 93.11 $\pm$ 0.36 3.70 $\pm$ 0.25 1.56 $\pm$ 0.18 92.81 $\pm$ 0.49 4.12 $\pm$ 0.25 1.61 $\pm$ 0.14
MDCA (Hebbalaguppe et al., 2022b) 93.05 $\pm$ 0.17 3.96 $\pm$ 0.12 1.63 $\pm$ 0.10 93.36 $\pm$ 0.24 4.30 $\pm$ 0.13 1.66 $\pm$ 0.10 92.03 $\pm$ 0.23 4.20 $\pm$ 0.16 1.63 $\pm$ 0.11
NUCFL (MDCA+COS) 92.59 $\pm$ 0.28 3.78 $\pm$ 0.30 1.60 $\pm$ 0.15 93.14 $\pm$ 0.24 4.04 $\pm$ 0.12 1.63 $\pm$ 0.06 92.02 $\pm$ 0.19 4.12 $\pm$ 0.22 1.60 $\pm$ 0.14
NUCFL (MDCA+L-CKA) 92.57 $\pm$ 0.21 3.70 $\pm$ 0.25 1.58 $\pm$ 0.13 93.12 $\pm$ 0.55 3.87 $\pm$ 0.41 1.60 $\pm$ 0.21 92.13 $\pm$ 0.24 4.10 $\pm$ 0.20 1.61 $\pm$ 0.13
NUCFL (MDCA+RBF-CKA) 92.65 $\pm$ 0.11 3.61 $\pm$ 0.12 1.53 $\pm$ 0.09 93.21 $\pm$ 0.17 4.12 $\pm$ 0.25 1.61 $\pm$ 0.11 92.14 $\pm$ 0.31 4.10 $\pm$ 0.14 1.60 $\pm$ 0.08
Calibration Method FedDyn (Acar et al., 2021) FedNova (Wang et al., 2020b)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 92.53 $\pm$ 0.25 5.13 $\pm$ 0.18 2.55 $\pm$ 0.09 92.91 $\pm$ 0.31 4.75 $\pm$ 0.17 2.38 $\pm$ 0.08
Focal (Lin et al., 2017) 91.63 $\pm$ 0.27 6.08 $\pm$ 0.24 3.03 $\pm$ 0.11 91.33 $\pm$ 0.42 4.75 $\pm$ 0.27 2.39 $\pm$ 0.19
LS (Müller et al., 2019) 91.84 $\pm$ 0.28 6.88 $\pm$ 0.17 3.48 $\pm$ 0.19 91.13 $\pm$ 0.37 5.40 $\pm$ 0.16 2.60 $\pm$ 0.11
BS (Brier, 1950) 91.40 $\pm$ 0.61 5.23 $\pm$ 0.21 2.57 $\pm$ 0.17 91.44 $\pm$ 0.71 4.81 $\pm$ 0.32 2.40 $\pm$ 0.19
MMCE (Kumar et al., 2018) 93.30 $\pm$ 0.49 7.73 $\pm$ 0.42 3.85 $\pm$ 0.14 92.44 $\pm$ 0.61 5.96 $\pm$ 0.68 3.03 $\pm$ 0.37
FLSD (Mukhoti et al., 2020) 92.40 $\pm$ 0.27 5.03 $\pm$ 0.16 2.32 $\pm$ 0.08 91.95 $\pm$ 0.40 4.94 $\pm$ 0.33 2.44 $\pm$ 0.19
MbLS (Liu et al., 2021) 91.60 $\pm$ 0.30 5.18 $\pm$ 0.32 2.56 $\pm$ 0.17 92.00 $\pm$ 0.55 4.88 $\pm$ 0.42 2.41 $\pm$ 0.17
DCA (Liang et al., 2020) 93.00 $\pm$ 0.27 4.60 $\pm$ 0.16 2.30 $\pm$ 0.13 93.06 $\pm$ 0.26 4.25 $\pm$ 0.17 1.68 $\pm$ 0.13
NUCFL (DCA+COS) 93.08 $\pm$ 0.35 4.38 $\pm$ 0.29 1.67 $\pm$ 0.17 93.26 $\pm$ 0.23 4.14 $\pm$ 0.30 1.59 $\pm$ 0.17
NUCFL (DCA+L-CKA) 92.87 $\pm$ 0.26 4.22 $\pm$ 0.27 1.64 $\pm$ 0.15 93.06 $\pm$ 0.24 4.03 $\pm$ 0.27 1.54 $\pm$ 0.11
NUCFL (DCA+RBF-CKA) 93.09 $\pm$ 0.27 4.25 $\pm$ 0.19 1.65 $\pm$ 0.11 92.99 $\pm$ 0.25 4.15 $\pm$ 0.17 1.58 $\pm$ 0.10
MDCA (Hebbalaguppe et al., 2022b) 92.70 $\pm$ 0.26 4.77 $\pm$ 0.14 2.38 $\pm$ 0.08 92.89 $\pm$ 0.25 4.37 $\pm$ 0.13 1.66 $\pm$ 0.09
NUCFL (MDCA+COS) 93.25 $\pm$ 0.19 4.41 $\pm$ 0.14 1.68 $\pm$ 0.09 93.06 $\pm$ 0.22 4.21 $\pm$ 0.22 1.64 $\pm$ 0.10
NUCFL (MDCA+L-CKA) 92.87 $\pm$ 0.26 4.29 $\pm$ 0.24 1.65 $\pm$ 0.11 92.85 $\pm$ 0.07 4.17 $\pm$ 0.09 1.62 $\pm$ 0.09
NUCFL (MDCA+RBF-CKA) 93.00 $\pm$ 0.15 4.36 $\pm$ 0.22 1.66 $\pm$ 0.10 92.92 $\pm$ 0.19 4.22 $\pm$ 0.13 1.64 $\pm$ 0.07
@@
Table 11: Accuracy (%), calibration measures ECE (%), and SCE (%) of various FL algorithms with different calibration methods under IID scenario on the FEMNIST dataset. Underlined values indicate the best calibration across all methods.
@@
Calibration Method FedAvg (McMahan et al., 2017) FedProx (Sahu et al., 2018) Scaffold (Karimireddy et al., 2019)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 90.95 $\pm$ 0.64 4.17 $\pm$ 1.04 1.66 $\pm$ 0.72 91.45 $\pm$ 1.17 4.61 $\pm$ 1.03 1.95 $\pm$ 0.99 91.28 $\pm$ 1.29 5.02 $\pm$ 1.19 2.51 $\pm$ 0.88
Focal (Lin et al., 2017) 90.63 $\pm$ 0.98 4.77 $\pm$ 1.30 1.96 $\pm$ 0.75 90.00 $\pm$ 1.29 5.39 $\pm$ 1.43 2.77 $\pm$ 1.05 91.13 $\pm$ 1.66 5.38 $\pm$ 1.27 2.72 $\pm$ 1.04
LS (Müller et al., 2019) 91.07 $\pm$ 1.23 3.75 $\pm$ 1.72 1.62 $\pm$ 1.02 90.37 $\pm$ 1.38 4.52 $\pm$ 1.14 1.90 $\pm$ 1.04 91.33 $\pm$ 1.75 5.18 $\pm$ 1.08 2.64 $\pm$ 0.99
BS (Brier, 1950) 91.48 $\pm$ 0.88 5.19 $\pm$ 1.24 2.65 $\pm$ 0.83 90.98 $\pm$ 1.00 5.42 $\pm$ 0.95 2.80 $\pm$ 0.61 88.72 $\pm$ 1.11 4.19 $\pm$ 1.43 1.81 $\pm$ 1.06
MMCE (Kumar et al., 2018) 90.22 $\pm$ 1.11 4.85 $\pm$ 1.23 2.30 $\pm$ 1.05 90.12 $\pm$ 1.32 4.01 $\pm$ 2.04 1.79 $\pm$ 1.38 91.44 $\pm$ 1.33 5.07 $\pm$ 1.68 2.55 $\pm$ 1.21
FLSD (Mukhoti et al., 2020) 90.02 $\pm$ 1.37 4.93 $\pm$ 1.28 2.95 $\pm$ 0.74 90.39 $\pm$ 1.25 4.99 $\pm$ 1.38 2.04 $\pm$ 1.06 92.88 $\pm$ 0.84 5.19 $\pm$ 1.14 2.58 $\pm$ 0.92
MbLS (Liu et al., 2021) 90.62 $\pm$ 1.59 4.27 $\pm$ 1.03 1.99 $\pm$ 0.80 91.49 $\pm$ 1.56 4.79 $\pm$ 1.50 1.94 $\pm$ 0.84 92.87 $\pm$ 1.64 5.39 $\pm$ 1.35 2.60 $\pm$ 0.91
DCA (Liang et al., 2020) 91.84 $\pm$ 1.06 3.61 $\pm$ 0.92 1.52 $\pm$ 0.53 92.03 $\pm$ 1.04 4.25 $\pm$ 0.58 1.61 $\pm$ 0.69 92.04 $\pm$ 1.30 4.44 $\pm$ 1.08 1.82 $\pm$ 0.88
NUCFL (DCA+COS) 91.77 $\pm$ 1.02 3.52 $\pm$ 0.79 1.49 $\pm$ 0.45 91.95 $\pm$ 1.28 3.61 $\pm$ 1.04 1.35 $\pm$ 0.96 92.42 $\pm$ 1.02 4.39 $\pm$ 1.37 1.77 $\pm$ 1.08
NUCFL (DCA+L-CKA) 91.63 $\pm$ 0.95 3.52 $\pm$ 0.88 1.47 $\pm$ 0.61 92.10 $\pm$ 0.96 3.60 $\pm$ 0.88 1.30 $\pm$ 0.85 92.19 $\pm$ 0.95 4.09 $\pm$ 0.83 1.60 $\pm$ 0.65
NUCFL (DCA+RBF-CKA) 91.74 $\pm$ 1.15 3.49 $\pm$ 1.00 1.40 $\pm$ 0.78 91.99 $\pm$ 1.23 3.77 $\pm$ 1.04 1.35 $\pm$ 1.03 91.74 $\pm$ 1.33 4.15 $\pm$ 1.29 1.62 $\pm$ 0.94
MDCA (Hebbalaguppe et al., 2022b) 91.64 $\pm$ 0.85 3.75 $\pm$ 1.29 1.53 $\pm$ 0.95 92.17 $\pm$ 1.33 4.42 $\pm$ 1.39 2.05 $\pm$ 1.06 92.95 $\pm$ 1.38 4.61 $\pm$ 1.29 1.90 $\pm$ 1.07
NUCFL (MDCA+COS) 91.29 $\pm$ 0.95 3.61 $\pm$ 1.20 1.44 $\pm$ 1.04 91.95 $\pm$ 1.00 3.95 $\pm$ 1.29 1.77 $\pm$ 0.93 93.07 $\pm$ 1.19 4.14 $\pm$ 1.22 1.62 $\pm$ 0.57
NUCFL (MDCA+L-CKA) 91.97 $\pm$ 1.06 3.28 $\pm$ 1.11 1.28 $\pm$ 0.93 92.20 $\pm$ 1.20 3.88 $\pm$ 1.04 1.51 $\pm$ 1.22 92.77 $\pm$ 1.33 4.20 $\pm$ 1.29 1.80 $\pm$ 1.08
NUCFL (MDCA+RBF-CKA) 91.35 $\pm$ 1.13 3.56 $\pm$ 0.99 1.40 $\pm$ 0.61 92.21 $\pm$ 0.95 4.00 $\pm$ 1.33 1.77 $\pm$ 1.16 93.04 $\pm$ 1.52 4.19 $\pm$ 1.63 1.79 $\pm$ 1.05
Calibration Method FedDyn (Acar et al., 2021) FedNova (Wang et al., 2020b)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 92.75 $\pm$ 1.56 5.24 $\pm$ 1.22 2.60 $\pm$ 1.03 92.22 $\pm$ 1.00 4.92 $\pm$ 0.95 2.44 $\pm$ 0.61
Focal (Lin et al., 2017) 91.15 $\pm$ 1.07 5.84 $\pm$ 1.05 2.91 $\pm$ 0.84 91.13 $\pm$ 1.20 5.39 $\pm$ 1.49 2.63 $\pm$ 0.92
LS (Müller et al., 2019) 93.08 $\pm$ 1.90 5.19 $\pm$ 1.65 2.63 $\pm$ 1.01 92.04 $\pm$ 1.11 4.95 $\pm$ 1.32 2.46 $\pm$ 0.85
BS (Brier, 1950) 91.39 $\pm$ 1.24 5.37 $\pm$ 0.98 2.62 $\pm$ 0.86 90.38 $\pm$ 1.33 4.65 $\pm$ 1.52 1.92 $\pm$ 1.04
MMCE (Kumar et al., 2018) 92.11 $\pm$ 1.43 4.93 $\pm$ 1.29 2.49 $\pm$ 1.00 91.35 $\pm$ 1.59 5.08 $\pm$ 1.29 2.49 $\pm$ 0.96
FLSD (Mukhoti et al., 2020) 92.06 $\pm$ 1.69 5.44 $\pm$ 1.77 2.75 $\pm$ 1.24 90.39 $\pm$ 1.71 4.61 $\pm$ 1.33 1.92 $\pm$ 1.06
MbLS (Liu et al., 2021) 92.06 $\pm$ 1.69 5.44 $\pm$ 1.77 2.75 $\pm$ 1.24 90.39 $\pm$ 1.71 4.61 $\pm$ 1.33 1.92 $\pm$ 1.06
DCA (Liang et al., 2020) 93.08 $\pm$ 1.44 4.61 $\pm$ 1.58 1.92 $\pm$ 1.06 92.37 $\pm$ 1.35 4.31 $\pm$ 1.24 1.71 $\pm$ 0.91
NUCFL (DCA+COS) 93.22 $\pm$ 1.11 4.20 $\pm$ 1.25 1.65 $\pm$ 0.93 92.25 $\pm$ 1.04 4.13 $\pm$ 0.95 1.68 $\pm$ 0.74
NUCFL (DCA+L-CKA) 93.45 $\pm$ 0.96 4.22 $\pm$ 1.14 1.65 $\pm$ 0.69 92.33 $\pm$ 1.22 4.02 $\pm$ 1.09 1.54 $\pm$ 0.83
NUCFL (DCA+RBF-CKA) 92.95 $\pm$ 1.61 4.34 $\pm$ 1.33 1.75 $\pm$ 1.10 92.41 $\pm$ 1.15 4.11 $\pm$ 1.03 1.68 $\pm$ 0.65
MDCA (Hebbalaguppe et al., 2022b) 93.19 $\pm$ 1.07 4.71 $\pm$ 0.94 1.93 $\pm$ 0.75 92.05 $\pm$ 1.28 4.62 $\pm$ 1.14 1.82 $\pm$ 0.85
NUCFL (MDCA+COS) 93.11 $\pm$ 1.64 4.31 $\pm$ 1.22 1.75 $\pm$ 0.99 91.99 $\pm$ 1.22 4.28 $\pm$ 0.98 1.67 $\pm$ 0.59
NUCFL (MDCA+L-CKA) 93.11 $\pm$ 1.64 4.31 $\pm$ 1.22 1.75 $\pm$ 0.99 91.99 $\pm$ 1.22 4.28 $\pm$ 0.98 1.67 $\pm$ 0.59
NUCFL (MDCA+RBF-CKA) 93.04 $\pm$ 1.23 4.40 $\pm$ 1.35 1.80 $\pm$ 1.05 92.39 $\pm$ 0.95 4.17 $\pm$ 0.93 1.67 $\pm$ 0.85
@@
Table 12: Accuracy (%), calibration measures ECE (%), and SCE (%) of various FL algorithms with different calibration methods under non-IID ($\alpha=0.5$) scenario on the FEMNIST dataset. Values in boldface represent the best calibration provided by our method for the auxiliary calibration method, and underlined values indicate the best calibration across all methods.